64 identical drops each of capacity 5microfarad combine to form a big drop. What is the capacity of big drop

Grade:12

2 Answers

Vikas TU
7 years ago
Capacitance for drop => R/k
where R is the radii.
From Volume conservation,
Vnet = 64V
R^3 = 64r^3
R = 4r
For big drop,
C = R/k.......(1)
For small dop,
5 = r/k........(2)
Divdng the eqns. we get,
C = 5*(R/r) => 5*4 = 20 mF.
 
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

We first find volume of big drop volume of big drop = 64 x volume of small drop
4/3 pi. R^3 = 64 x 4/3 pi. r^3
R = 4r
because capacitance of solid spherical body directly proportion to radius
hence big drop capacitance = 4 x capacitance of small
capacitance of big drop = 20 microfarad

Thanks and Regards